Fermi National Laboratory
 

Accelerator Update for October 15, 2007 - October 16, 2007

 

Monday October 15

The day shift began with the accelerator complex in startup.

Crew Chief Summary: Experts in the Booster, Main Injector (MI), the Antiproton Source (Pbar), and Recycler conducted beam studies. MiniBooNE/SciBooNE experts did some beam tuning. The beam line was then set to run in autotune mode and ran that way overnight. Tevatron (TeV) sectors F4 cryo system problems were fixed late on the day shift. The TeV cooldown was complete by 7:56 PM. Permission was granted to power the low beta magnet power supplies. Permission was granted to run beam in the Tevatron. TeV experts started the TeV turn on. Operators replaced several DFG regulators in the TeV. Problems with AQ7, DQ0, DQ7, and DQ9 prevented a successful low beta checkout. A QBS ramp was established on the midnight shift. EE Support arrived at 6 AM to begin working on the new dump/bypass hardware in the TeV. Experts tuned the Linac and Booster overnight. Operators prepared the Booster, MI, and Pbar Transport areas for controlled accesses.

Tuesday October 16

The midnight shift began with the accelerator complex in startup.

Crew Chief Summary: Operators prepared the Pbar rings for controlled access on the day shift. Experts and techs accessed the Booster, MI, Pbar Rings and Transport for scheduled work. EE Support began TeV power supply studies. MiniBooNE reported that they were ready for beam when it becomes available.

Machine Summaries

Linac: Delivered beam to Booster. Experts were working on an RF station (KRF1) that had an instability problem. Experts were also working on low-level-RF trigger problems and their velocity meter.

Booster: Delivered beam to the Main Injector. Experts were tuning the beam.

Antiproton Source: Experts had established beam to the Accumulator and the Pbar Debuncher, but not to AP2 yet. They were having problems with some application programs. They found a left bend magnet (IB7) with hot coils, which will need to be back-flushed. Experts need to work on shunts, and to measure orbits and apertures. They want to begin stacking and phase cooling.

Main Injector: Delivered beam to the Antiproton Source. Technicians found an LCW leak in the 303 region that needs re-brazing. They have narrowed a ground fault search to near the collimators; it may be due to a thermal problem; experts will change the turn on order. If that doesn’t work they will have to access and remove shielding to investigate further. Experts need to set extraction positions.

Recycler: Their experts are waiting for antiprotons. They can use protons for studies. They will need at least one day of studies with antiprotons before they can send antiprotons to the TeV.

Tevatron: The cryo system went cold at 8 PM on Monday night. Operations turned on to a QBS ramp. They found that all of their pond water pumps had been turned off. Experts reported that a wet engine needed a valve change and that a controller crate power supply connector needed changing. The experts want to inject protons into the TeV this evening.

SwitchYard 120: The first user will have their experiment in place and want beam by November 7 th. Switchyard experts will need beam at least a week before that for tuning.

NuMI: Technicians continue to work on their humidifiers and evaporators. They are waiting on instrumentation cable pulls. Experts determined that their Horn #2 ground fault was due to metal flakes.

MiniBooNE/SciBooNE: Experts reported that they are ready for beam.

D0: Their experts reported that they found a bad power supply fuse that will require them to open up their detector, which will take an 8-hour access to repair. They continued to take cosmic rays.

CDF: Their experts reported that they have been working on wire chambers and taking cosmic rays.

FESS: They reported that they are replacing a suction valve for TeV sector D3. The Pbar pump #8 motor went bad; they switched on pump #7. Experts reported that they have started the boilers in preparation for the heating season.

Mechanical: Experts were flushing a Pbar magnet and repairing LCW leaks.

The Plans for Tuesday

The plans for today are to access the Booster, MI, and the Pbar rings and transport areas for necessary work.


Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 10/16/2007    email Fermilab